A retiree receives $7500 a year interest from $80,000 placed in two bonds, one paying 6% interest and the other paying 16% interest. How much is being invested at the higher interest rate.
Answer by lwsshak3(11628)

A retiree receives $7500 a year interest from $80,000 placed in two bonds, one paying 6% interest and the other paying 16% interest. How much is being invested at the higher interest rate.
***
let x=amt invested at 16% interest
80000-x=amt invested at 6% interest
..
16%x+6%(80000-x)=7500
.16x+4800-.06x=7500
.10x=2700
x=27000
How much is being invested at the higher interest rate? $27,000